Solve by elimination: * 7x + 2y = 24 / 8x + 2y = 30
Answer by Cromlix(4381)

7x + 2y = 24
8x + 2y = 30
Subtract
-x = -6
x = 6
Substitute x = 6
into 7x + 2y = 24
7(6)+ 2y = 24
42 + 2y = 24
2y = 24 - 42
2y = -18
y = -9
{6,-9}